The tax rates are as shown.

Taxable Income Tax Rate
$0 – 50,000 15%
50,001 – 75,000 25%
75,001 – 100,000 34%
100,001 – 335,000 39%

    

Your firm currently has taxable income of $80,000. How much additional tax will you owe if you increase your taxable income by $21,200?
Answer #1

Taxes when the income is $80000

= 50000 * 15% + (75000 - 50000) * 25% + (80000 - 75000) *34%

= 15450

Taxes when income is (80000 + 21200) = 101200

= 50000 * 15% + (75000 - 50000) * 25% + (100000 - 75000) *34% + (101200 - 100000) *39%

= 22718

Additional taxes owed = 22718 - 15450 = 7268